# On Creating a Causally Grounded Usable Rating Method for Assessing the Robustness of Foundation Models Supporting Time Series

### v1

@@ 4 Solution Approach
Perturbations: We introduce one syntactic (STP) and two semantic perturbations (SMP) (Figure 3(a)) inspired by real-world applications in unintended scenarios to assess the accuracy and robustness of FMTS. SMPs alter data meaning while preserving context, e.g., a stock’s value might change due to data entry errors or market-specific catalysts. STPs modify the structure of the data while preserving without altering the content. Drop-to-zero (P1) is an SMP inspired by common data entry errors (Ley et al. 2019). Every $n^{th}$ value in the original stock price data is set to zero. Sampling the time series with a sliding window of size $n$ ensures each sample contains a zero. In Value halved perturbation (P2), an SMP, we reduce every $n^{th}$ number in the original stock price data to half of its value. This perturbation simulates periodic adjustments, possibly reflecting events like stock splits or dividend payments. Missing values perturbation (P3), an STP, converts every $n^{th}$ number in the original stock price data to a null value, simulating real-world missing data points in financial datasets due to system incomplete transmissions.
@@ 5.2 Experimental Evaluation
Figure 5: Radar plots showing for all systems (a) mean forecasting accuracy with respect to all metrics, (b) forecasting accuracy under P2 (half-valued pertubation) Appendix E: Figs. 18, 19 show all perturbations; (c) mean robustness metrics for FMTS and $S_{a}$, and (d) robustness under P2. Each axis is normalized and inverted if needed so that outer ring implies better performance.
RQ1: Does Sensitive Attribute affect the Residual, even though Sensitive Attribute has no effect on Perturbation? Setup: In this experiment, the causal link from the Sensitive Attribute to Perturbation is absent, as the perturbation to the stock prices does not depend on the corresponding company name or the industry i.e., perturbations are applied uniformly across all the data points. We quantify the statistical bias exhibited by the systems by using WRS described in Section 4.1. We perform two different analyses in this experiment: one to measure the discrepancy shown across different industries (WRSIndustry) and another to measure the discrepancy among both the companies (WRSCompany) within the same industry. Results and conclusion: From Figure 4, most discrepancies can be observed across industries compared to the discrepancies across companies within each industry. When the input data was subjected to perturbation P3 ($WRS_{avg}$ of 5.35333Table 4 in Appendix E shows the RQs along with the average values for all metrics across different systems and perturbations.), the systems exhibited more statistical bias. From Figure 5, $S_{a}$ ($WRS_{avg}$ of 3.96) exhibited the least statistical bias, while $S_{p}$ ($WRS_{avg}$ of 6.28) exhibited the highest statistical bias among the systems evaluated under the perturbations considered. Hence, we conclude that Sensitive Attribute affects the Residual, even though Sensitive Attribute has no effect on Perturbation.
@@
Setup: In this experiment, we use PIE % defined in equation 2 to compare the APE (defined in equation 1) before and after deconfounding using the PSM technique as the presence of the confounder opens a backdoor path from Perturbation to Residual through the Confounder. The causal link from Confounder to Perturbation will be valid only if the perturbation applied depends on the value of the confounder (i.e. the company or the industry the specific data points belong to). To ensure the probability of perturbation assignment varies with respect to the Confounder across three distributions (DI1 through DI3) in the case of Industry and six different distributions in the case of Company (DC1 through DC6), we implement weighted sampling. For each distribution, weights are configured so that perturbation groups P1 through P3 have a twofold higher likelihood of selection compared to P0 for specific values of the confounder. For example, META in DC1, GOOG in DC2, and so on. This strategy highlights significant cases, although other combinations are possible for further exploration. Results and Conclusion: Figure 4 shows that selecting Company as the confounder leads to greater confounding bias in the systems. In Figure 5, $S_{g}^{ni}$ ($PIE_{avg}\%$ of 523.09) exhibited the least confounding bias, while $S_{p}$ ($PIE_{avg}\%$ of 2677.62) exhibited the most. Systems showed more confounding bias under perturbation P3 ($PIE_{avg}\%$ of 1563.94). Therefore, the Confounder affects the relationship between Perturbation and Residual, particularly when the Confounder influences the Perturbation.
@@
RQ3: Does Perturbation affect the Residual when Sensitive Attribute may have an effect on Residual? Setup: The experimental setup in this experiment is same as that for answering RQ2. To compute the APE, we used PSM described in Section 4.1. PSM allows us to effectively determine the effect of Perturbation on the Residual. For instance, if two matched points belong to the same company but only one was perturbed, any difference in their residuals can be directly attributed to the perturbation itself rather than to other confounding factors. This method provides a clear understanding of the true impact of the Perturbation on the Residual. As our rating method aims to bring the worst possible behavior of the systems, we take the MAX(APE) as the raw score that is used to compute the final ratings. Results and Conclusion: It is undesirable to have a higher APE, as it implies that the perturbation applied can have a significant impact on the residuals of different systems. From Figure 4, when Industry was considered as the confounder, it led to a higher APE. As the outcome of $S_{b}$ depended on the Company (and varied from one company to another), the perturbation did not have any effect on the system. Whereas, when Industry was considered as the confounder, the perturbation appeared influential, resulting in a high APE for $S_{b}$. From Figure 5, perturbations had the least impact on $S_{g}^{ni}$ ($APE_{avg}$ of 6.49) and highest impact on $S_{p}$ ($APE_{avg}$ of 28.53). Among all the perturbations, P1 ($APE_{avg}$ of 21.11) was the most disruptive. Hence, Perturbation affects the Residual when Sensitive Attribute may have an effect on Residual.
@@
RQ4: Does Perturbations degrade the accuracy of S? Setup: For this experiment, we do not use any causal model. We compute the three accuracy metrics widely used in for the task of financial time-series forecasting (Makridakis, Spiliotis, and Assimakopoulos 2022) which were summarized in Section 4.1. Results and Conclusion: From Figure 5, $S_{c}$ exhibited the highest amount of forecasting accuracy in terms of SMAPE (average of 0.05) and MASE (average of 4.67), while $S_{a}$ outperformed all other systems in terms of sign accuracy (average of 58.57). $S_{b}$ consistently predicted the correct directional movement of stock prices, exhibiting high sign accuracy as it was designed to adjust residuals based on specific company stock prices. Perturbations P2, P1, and P0 caused the highest decline in SMAPE (average of 0.08), MASE (average of 8.99), and sign accuracy (average of 49.97), respectively. Hence, Perturbations degrade the accuracy of S.
@@ 5.3 Overall Performance Comparison
Now, we provide an overall comparison of the different systems across all metrics to highlight key findings about their performance under various perturbations. Figure 5 shows radar plots with mean forecasting accuracy and robustness metrics, including values under perturbation P2.
1. Clear Domination Signals From Figure 5 and the detailed results from Section 5.2, we can draw the following conclusions: $S_{c}$’s Superiority in Forecasting Metrics: $S_{c}$ consistently outperformed other models in terms of forecasting accuracy metrics, specifically SMAPE and MASE. This indicates that $S_{c}$ is highly effective in predicting stock prices with minimal error. General Superiority Over Biased and Random Systems: All models perform better than the biased and random systems in forecasting metrics. This underscores the importance of using well-designed models over naive or biased approaches. Robustness in PIE % and APE Metrics: According to the average scores, the $S_{g}^{ni}$ system demonstrated superior robustness in PIE % and APE metrics. This suggests that $S_{g}^{ni}$ is more resilient to perturbations and confounding biases compared to other systems.
@@
3. Role of Confounders Our analysis (Figs. 4 and 5: left) shows that using industry as a confounder introduces more bias, with higher PIEscores indicating significant industry-specific effects on the relationship between perturbations and residuals. Inter-industry comparisons also show more discrepancies, as evidenced by WRS scores. Conversely, the impact of company as a confounder varies. For instance, system $S_{b}$ shows minimal effect from company-specific perturbations, suggesting it is well-tailored to company data, while other systems have higher APE scores, indicating a significant impact on residuals due to company-specific factors.
@@ Appendix D Figures in a Higher Resolution from Main Paper
Figure 13: Radar plots showing for all systems (a) mean forecasting accuracy with respect to all metrics, (b) forecasting accuracy under P2 (half-valued pertubation) Appendix E: Figs. 18, 19 show all perturbations; (c) mean robustness metrics for FMTS and $S_{a}$, and (d) robustness under P2. Each axis is normalized and inverted if needed so that outer ring implies better performance.
@@ Appendix E Additional Experimental Results
In this section, Table 2 shows the partial order and final order with respect to all the metrics defined in Section 4.1. Table 3 shows the forecasting accuracy values for all the systems. Table 4 shows the research questions, and average values for all metrics across systems and perturbations (average values are referred to in the conclusions made for each RQ in Section 5.2). Figure 16 and 17 shows bar plots with all the robustness metric values and forecasting accuracy values. Figure 20 shows the heatmap for all metrics for all the models.
Table 2 shows the partial order and final order with respect to all the metrics defined in Section 4.1. Table 3 shows the forecasting accuracy values for all the systems. Table 4 shows the research questions, and average values for all metrics across systems and perturbations (average values are referred to in the conclusions made for each RQ in Section 5.2). Figure 16 and 17 shows bar plots with all the robustness metric values and forecasting accuracy values. Figure 20 shows the heatmap for all metrics for all the models.

### v2

@@ 4 Solution Approach
Perturbations: We introduce one syntactic (STP) and two semantic perturbations (SMP) (Figure 3(a)) inspired by real-world applications in unintended scenarios to assess the accuracy and robustness of FMTS. SMPs alter data meaning while preserving context, e.g., a stock’s value might change due to data entry errors or market-specific catalysts. STPs modify the structure of the data without altering the content. Drop-to-zero (P1) is an SMP inspired by common data entry errors (Ley et al. 2019). Every $n^{th}$ value in the original stock price data is set to zero. Sampling the time series with a sliding window of size $n$ ensures each sample contains a zero. In Value halved perturbation (P2), an SMP, we reduce every $n^{th}$ number in the original stock price data to half of its value. This perturbation simulates periodic adjustments, possibly reflecting events like stock splits or dividend payments. Missing values perturbation (P3), an STP, converts every $n^{th}$ number in the original stock price data to a null value, simulating real-world missing data points in financial datasets due to system incomplete transmissions.
@@ 5.2 Experimental Evaluation
RQ1: Does Sensitive Attribute affect the Residual, even though Sensitive Attribute has no effect on Perturbation? Setup: In this experiment, the causal link from the Sensitive Attribute to Perturbation is absent, as the perturbation to the stock prices does not depend on the corresponding company name or the industry i.e., perturbations are applied uniformly across all the data points. We quantify the statistical bias exhibited by the systems by using WRS described in Section 4.1. We perform two different analyses in this experiment: one to measure the discrepancy shown across different industries (WRSIndustry) and another to measure the discrepancy among both the companies (WRSCompany) within the same industry. Results and conclusion: From Figure 4, most discrepancies can be observed across industries compared to the discrepancies across companies within each industry. When the input data was subjected to perturbation P3 ($WRS_{avg}$ of 5.35333Table 4 in Appendix E shows the RQs along with the average values for all metrics across different systems and perturbations.), the systems exhibited more statistical bias. From Figure LABEL:fig:radar-agg, $S_{a}$ ($WRS_{avg}$ of 3.96) exhibited the least statistical bias, while $S_{p}$ ($WRS_{avg}$ of 6.28) exhibited the highest statistical bias among the systems evaluated under the perturbations considered. Hence, we conclude that Sensitive Attribute affects the Residual, even though Sensitive Attribute has no effect on Perturbation.
@@
Setup: In this experiment, we use PIE % defined in equation 2 to compare the APE (defined in equation 1) before and after deconfounding using the PSM technique as the presence of the confounder opens a backdoor path from Perturbation to Residual through the Confounder. The causal link from Confounder to Perturbation will be valid only if the perturbation applied depends on the value of the confounder (i.e. the company or the industry the specific data points belong to). To ensure the probability of perturbation assignment varies with respect to the Confounder across three distributions (DI1 through DI3) in the case of Industry and six different distributions in the case of Company (DC1 through DC6), we implement weighted sampling. For each distribution, weights are configured so that perturbation groups P1 through P3 have a twofold higher likelihood of selection compared to P0 for specific values of the confounder. For example, META in DC1, GOOG in DC2, and so on. This strategy highlights significant cases, although other combinations are possible for further exploration. Results and Conclusion: Figure 4 shows that selecting Company as the confounder leads to greater confounding bias in the systems. In Figure LABEL:fig:radar-agg, $S_{g}^{ni}$ ($PIE_{avg}\%$ of 523.09) exhibited the least confounding bias, while $S_{p}$ ($PIE_{avg}\%$ of 2677.62) exhibited the most. Systems showed more confounding bias under perturbation P3 ($PIE_{avg}\%$ of 1563.94). Therefore, the Confounder affects the relationship between Perturbation and Residual, particularly when the Confounder influences the Perturbation.
@@
RQ3: Does Perturbation affect the Residual when Sensitive Attribute may have an effect on Residual? Setup: The experimental setup in this experiment is same as that for answering RQ2. To compute the APE, we used PSM described in Section 4.1. PSM allows us to effectively determine the effect of Perturbation on the Residual. For instance, if two matched points belong to the same company but only one was perturbed, any difference in their residuals can be directly attributed to the perturbation itself rather than to other confounding factors. This method provides a clear understanding of the true impact of the Perturbation on the Residual. As our rating method aims to bring the worst possible behavior of the systems, we take the MAX(APE) as the raw score that is used to compute the final ratings. Results and Conclusion: It is undesirable to have a higher APE, as it implies that the perturbation applied can have a significant impact on the residuals of different systems. From Figure 4, when Industry was considered as the confounder, it led to a higher APE. As the outcome of $S_{b}$ depended on the Company (and varied from one company to another), the perturbation did not have any effect on the system. Whereas, when Industry was considered as the confounder, the perturbation appeared influential, resulting in a high APE for $S_{b}$. From Figure LABEL:fig:radar-agg, perturbations had the least impact on $S_{g}^{ni}$ ($APE_{avg}$ of 6.49) and highest impact on $S_{p}$ ($APE_{avg}$ of 28.53). Among all the perturbations, P1 ($APE_{avg}$ of 21.11) was the most disruptive. Hence, Perturbation affects the Residual when Sensitive Attribute may have an effect on Residual.
@@
RQ4: Does Perturbations degrade the accuracy of S? Setup: For this experiment, we do not use any causal model. We compute the three accuracy metrics widely used in for the task of financial time-series forecasting (Makridakis, Spiliotis, and Assimakopoulos 2022) which were summarized in Section 4.1. Results and Conclusion: From Figure LABEL:fig:radar-agg, $S_{c}$ exhibited the highest amount of forecasting accuracy in terms of SMAPE (average of 0.05) and MASE (average of 4.67), while $S_{a}$ outperformed all other systems in terms of sign accuracy (average of 58.57). $S_{b}$ consistently predicted the correct directional movement of stock prices, exhibiting high sign accuracy as it was designed to adjust residuals based on specific company stock prices. Perturbations P2, P1, and P0 caused the highest decline in SMAPE (average of 0.08), MASE (average of 8.99), and sign accuracy (average of 49.97), respectively. Hence, Perturbations degrade the accuracy of S.
@@ 5.3 Overall Performance Comparison
Now, we provide an overall comparison of the different systems across all metrics to highlight key findings about their performance under various perturbations. Figure LABEL:fig:radar-agg shows radar plots with mean forecasting accuracy and robustness metrics, including values under perturbation P2.
1. Clear Domination Signals From Figure LABEL:fig:radar-agg and the detailed results from Section 5.2, we can draw the following conclusions: $S_{c}$’s Superiority in Forecasting Metrics: $S_{c}$ consistently outperformed other models in terms of forecasting accuracy metrics, specifically SMAPE and MASE. This indicates that $S_{c}$ is highly effective in predicting stock prices with minimal error. General Superiority Over Biased and Random Systems: All models perform better than the biased and random systems in forecasting metrics. This underscores the importance of using well-designed models over naive or biased approaches. Robustness in PIE % and APE Metrics: According to the average scores, the $S_{g}^{ni}$ system demonstrated superior robustness in PIE % and APE metrics. This suggests that $S_{g}^{ni}$ is more resilient to perturbations and confounding biases compared to other systems.
@@
3. Role of Confounders Our analysis (Figs. 4 and LABEL:fig:radar-agg: left) shows that using industry as a confounder introduces more bias, with higher PIEscores indicating significant industry-specific effects on the relationship between perturbations and residuals. Inter-industry comparisons also show more discrepancies, as evidenced by WRS scores. Conversely, the impact of company as a confounder varies. For instance, system $S_{b}$ shows minimal effect from company-specific perturbations, suggesting it is well-tailored to company data, while other systems have higher APE scores, indicating a significant impact on residuals due to company-specific factors.
@@ 6 Discussion and Conclusion
Acknowledgements. This paper was prepared for informational purposes in part by the Artificial Intelligence Research group of JPMorgan Chase & Co and its affiliates (“J.P. Morgan”) and is not a product of the Research Department of J.P. Morgan.  J.P. Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein.  This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
@@ Appendix D Figures in a Higher Resolution from Main Paper
Figure 12: Radar plots showing for all systems (a) mean forecasting accuracy with respect to all metrics, (b) forecasting accuracy under P2 (half-valued pertubation) Appendix E: Figs. 17, 18 show all perturbations; (c) mean robustness metrics for FMTS and $S_{a}$, and (d) robustness under P2. Each axis is normalized and inverted if needed so that outer ring implies better performance.
@@ Appendix E Additional Experimental Results
In this section, Table 2 shows the partial order and final order with respect to all the metrics defined in Section 4.1. Table 3 shows the forecasting accuracy values for all the systems. Table 4 shows the research questions, and average values for all metrics across systems and perturbations (average values are referred to in the conclusions made for each RQ in Section 5.2). Figure 15 and 16 shows bar plots with all the robustness metric values and forecasting accuracy values. Figure 19 shows the heatmap for all metrics for all the models.
Table 2 shows the partial order and final order with respect to all the metrics defined in Section 4.1. Table 3 shows the forecasting accuracy values for all the systems. Table 4 shows the research questions, and average values for all metrics across systems and perturbations (average values are referred to in the conclusions made for each RQ in Section 5.2). Figure 15 and 16 shows bar plots with all the robustness metric values and forecasting accuracy values. Figure 19 shows the heatmap for all metrics for all the models.